# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 2. Preliminaries
In this section, we recall some basic concepts and results that will be used throughout this article, without proofs. For background on hyperbolic geometry concerning the hyperbolic plane and hyperbolic surfaces, we refer to Katok [Kat92], and for details on the Laplace–Beltrami operator, to Iwaniec [Iwa02].
2.1. The hyperbolic plane and hyperbolic surfaces
@@
and we write $|A|$ for the hyperbolic volume of a set $A\subseteq\mathbb{H}$.
The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm I$, acts on $\mathbb{H}$ by Möbius transformations:
@@
The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via
@@
Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$; hence, these two spaces can be identified.
@@
We denote by $z,z^{\prime}]$ the geodesic segment connecting two distinct points $z,z^{\prime}\in\mathbb{H}$, by $z,z^{\prime}](w)$ the tangent direction at a point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ its length.
@@
Then a point $x\in\mathbb{H}$ can be written as
@@
where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of pair in $T^{1}\mathbb{H}$.
The Laplacian on $\mathbb{H}$ is the differential operator
@@
which commutes with all isometries of $\mathbb{H}$.
@@ 2.2. Hyperbolic surfaces
where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$. In this way, $X$ can be represented by a fundamental domain in $\mathbb{H}$.
A function on $X$ can equivalently be regarded as a $\Gamma$-periodic function on $\mathbb{H}$, and vice versa. Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain.
@@
$\operatorname{InjRad}_{X}:=\sup_{z\in X}\operatorname{InjRad}_{X}(z).$
Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally to a Laplacian on the quotient $X=\Gamma\setminus\mathbb{H}$. To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on all smooth functions on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions. On a compact connected hyperbolic surface, the Laplacian $\Delta_{X}$ has a discrete spectrum
@@
3. Strategy of proofs
Three main theorems, Theorems 1.4, 1.5, and 1.7, are proven in Section 8. In this section we give a rough, overarching idea of the approach taken in the proofs. Most of the paper is dedicated for trying to establish Theorem 1.5. Proving this theorem is technical so the proof is split into multiple sections. These sections, Sections 4, 5, 6, and 7, are presented before the Section 8 so that the auxiliary results have been covered when the main theorems are under focus.
In Theorem 1.5 we want to bound the double sum appearing in Expression (3.3) for a fixed surface $X=\Gamma\setminus\mathbb{H}$ under suitable conditions. Hence we want to bound the following:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}$ (3.1)
We can split the inner product as follows:
$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.2)
when
@@ 3. Strategy of proofs
This split could be referred to as geometric-spectral split, since the denumerator contains the spectral data of the problem and the numerator the geometric data of the problem. $P_{t}$ is the hyperbolic wave propagator (see [LP82]). It is the quantization of the geodesic flow, hence a very natural object. Using this propagator makes the spectralside easy but complicates the analysis of the geometric side when comparing the setting to that of Le Masson and Sahlsten (see [LS17]) who used a ball averaging propagator which is not as natural as the wave propagator. $P_{t,\tau}$, a modified version of $P_{t}$, is used simultaneously with $P_{t}$ so that we can access the off-diagonal elements. This adds another layer of complexity to the proof when comparing to the setting in which only the diagonal elements are considered.
@@
The denumerator of Expression (3.2) contains the spectral data of the problem. We acquire bounds for the denominator in Section 4. Using the bounds acquired there gives an upper bound for Expression (3.1):
$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}aP_{t}\,dt\right\|_{HS}^{2},$
where the Hilbert-Schmidt norm upper bound was acquired by enlargening the sums to include all the basis elements after the spectral bound was pulled out of the sum. Now the Hilbert-Schmidt norm contains the geometric data of the problem.
Bouding the Hilbert-Schmidt norm is very technical. In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$. Using the acquired formulas, we bound the Hilbert-Schmidt norm in Section 6. This is based on carefully bounding the Hilbert-Schmidt norm expression and finally splitting it into two cases. In one we consider an integral over points with small injectivity radius and in the other an integral over points with large injectivity radius. To bound these two cases, we especially need the exponential mixing theorem, Theorem 1.2, and bounds for certain geometric weight $F_{t,t^{\prime},\rho}$ which appears in both the small and large injectivity radius cases, containing the core geometric information of the problem. Bounding $F_{t,t^{\prime},\rho}$ is done in Section 7. We remark that the exponential mixing theorem replaces the Nevo’s theorem appearing in the same role in [LS17] as the exponential mixing here.
With the Hilbert-Schmidt norm bounded, we acquire the final bounds for Expression (3.1) to conclude Theorem 1.5:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq E\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where we refer to Theorem 1.5 for the quantities appearing in the upper bound.
The limit version of the main theorems, Theorem 1.4, is proven using Theorem 1.5 together with the large-scale Weyl law of Le Masson and Sahlsten, Theorem 8.1. In the proof we want to prove that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have that
@@
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}\right)<\epsilon.$ (3.3)
Proving this proves only conditions (ii) and (iii); Le Masson and Sahlsten’s proof of quantum ergodicity covers case (i), see [LS17]. We remark here that the techniques used here can also be used to give a new proof for the quantum ergodicity theorem. Using the bounds acquired for the double sum in Theorem 1.5 gives the following upper bound for the previous expression:
@@
$\limsup_{n\to\infty}\left(\frac{1}{N(X_{n},I)}\,E\,\max(I)^{4}\,\frac{1}{T\,\beta(\lambda_{1}(X_{n}))^{3}}\,\left(\|a_{n}\|_{2}^{2}\,+\,8\pi^{2}\|a_{n}\|_{\infty}^{2}\,|D_{n}\setminus D_{n}(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X_{n}}}\right)\right),$
where we refer to Theorem 1.4 for the quantities appearing in the upper bound. Using property (BSC) helps us to get a universal lower bound for $\beta(\lambda_{1}(X_{n}))$ and using property (UND) helps us to get a universal lower bound for $\operatorname{InjRad}_{X_{n}}$. Weyl law together with property (BSC) shows that
$\frac{|D_{n}\setminus D_{n}(4T)|}{|N(X_{n},I)|}\xrightarrow{n\to\infty}0,$
as the Weyl law relates the growth of the number of eigenvalues in the energy window $I$ to that of the volume of the surface, $|X_{n}|$, and the ratio of the points with small injectivity radius and the volume of the surface goes to zero due to property (BSC). The core idea is to use properties (BSC), (EXP), and (UND) together with Weyl law to remove the limit supremum from the upper bound. After this, we have an upper bound in which $\delta$ appears. We can choose a suitable $\delta$ to get $\epsilon$ as an upper bound to conclude the proof.
To finally prove the probabilistic main theorem, Theorem 1.7, we take a very similar approach as with the proof of Theorem 1.4 as these theorems are very similar. We only prove Properties (ii) and (iii) of Theorem 1.4 as Le Masson and Sahlsten have already proved Property (i), see [LS24]. We remark that in probabilistic setting we do not impose Properties (BSC), (EXP), and (UND) and hence we do not have the large-scale Weyl law at out disposal. To remedy this, we get similar properties from the theory of random surfaces. These are recorded in Theorem 8.2. We can use these properties in the place of Properties (BSC), (EXP), and (UND) in a very similar way to get the conclusion to hold with high probability as we proceed in the sequence.
To conclude the section, we remark that the main novelty of this paper method-wise is the use of the exponential mixing theorem, Theorem 1.2, instead of using Nevos’s theorem to get the needed extra decay to establish Theorem 1.5. Another novelty is to separate a weight arising from the choice of propagator, denoted here by $F_{t,t^{\prime},\rho}$, which gives much required nimbleness in the proof; it can also be approximated separately without the observable function $a$ interfering.

### v2

@@ 2. Preliminaries
In this section, we recall some basic concepts and results that will be used throughout this article without proofs. For background on hyperbolic geometry concerning the hyperbolic plane and hyperbolic surfaces, we refer to Katok [KAT92], and for details on the Laplace–Beltrami operator, to Iwaniec [IWA02].
2.1. Hyperbolic plane
@@
and we denote by $|A|$ the hyperbolic volume of a set $A\subset\mathbb{H}$.
The group $PSL_{2}(\mathbb{R})$, consisting of $2\times 2$ real matrices with determinant one, modulo $\pm\mathrm{id}$, acts on $\mathbb{H}$ by Möbius transformations:
@@
The group $PSL_{2}(\mathbb{R})$ acts on $T^{1}\mathbb{H}$ via the action
@@
Evaluating this action at the point $(i,i)$ shows that $PSL_{2}(\mathbb{R})$ is homeomorphic to $T^{1}\mathbb{H}$. Thus these two spaces can be identified.
@@
Point $x\in\mathbb{H}$ can be written as
@@
where the polar coordinates are around $z$; $r$ and $\theta$ depend on $x$, and $\pi_{1}$ is the projection onto the first element of the pair in $T^{1}\mathbb{H}$.
We denote by $z,z^{\prime}]$ the geodesic segment going from $z\in\mathbb{H}$ to another point $z^{\prime}\in\mathbb{H}$. By $z,z^{\prime}](w)$ we denote the direction of the tangent vector of the geodesic at point $w\in[z,z^{\prime}]$, and by $|[z,z^{\prime}]|$ we denote the length of the geodesic segment.
The Laplace–Beltrami operator on $\mathbb{H}$ is the differential operator
@@
which commutes with all isometries of $\mathbb{H}$. We refer to it as the Laplacian operator.
@@ 2.2. Hyperbolic surfaces
where $\Gamma\subset PSL_{2}(\mathbb{R})$ is a discrete subgroup consisting solely of hyperbolic elements, that is, matrices whose trace satisfies $|\mathrm{tr}(\gamma)|>2$; $\Gamma$ is a Fuchsian group. In this way, $X$ can be represented by a compact subset of $\mathbb{H}$ called a fundamental domain.
A function on $X$ can equivalently be regarded as a $\Gamma$-periodic function on $\mathbb{H}$, and vice versa. Moreover, the integral of a function $f$ over $X$ equals the integral of the corresponding $\Gamma$-invariant function over a fundamental domain of $X$.
@@
$\operatorname{InjRad}_{X}:=\inf_{z\in X}\operatorname{InjRad}_{X}(z).$
Since the Laplacian on $\mathbb{H}$ commutes with isometries, it descends naturally on the quotient $X=\Gamma\setminus\mathbb{H}$. To distinguish the two, we denote by $\Delta_{\mathbb{H}}$ the Laplacian acting on $\mathbb{H}$, and by $\Delta_{X}$ the Laplacian acting on $\Gamma$-invariant functions. On a compact connected hyperbolic surface, the Laplacian $\Delta_{X}$ has a discrete spectrum
@@
The spectral gap of $\Delta_{X}$ is the difference between its two smallest eigenvalues.
3. Proof ideas
Three main theorems, Theorems 1.3, 1.4, and 1.5, are proven in Section 8 using the auxiliary results established in Sections 4, 5, 6, and 7. In this section we give an overarching idea of all the proofs to provide a coherent narrative; we refer to the latter sections for the exact definitions of the appearing quantities. We remark that in the actual proof, the coefficients are written carefully to make all dependencies explicit; here, we adopt a less explicit presentation.
The goal is to prove Properties (ii) and (iii) of Theorems 1.3 and 1.4. We prove the properties simultaneously: We want to show that for any $\epsilon>0$ there exists $\delta(\epsilon)>0$ such that for any $\tau\in\mathbb{R}$ we have
$\limsup_{n\to\infty}\frac{1}{N(X_{n},I)}\,\sum_{j:\rho_{j}^{(n)}\in I}\sum_{\begin{subarray}{c}k:\rho_{k}^{(n)}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta(\epsilon)\end{subarray}}|\langle\psi_{j}^{(n)},a_{n}\psi_{k}^{(n)}\rangle|^{2}<\epsilon.$ (3.1)
The setting is deterministic in Theorem 1.3 and probabilistic in Theorem 1.4. To achieve this goal, we aim to bound the double sum for a fixed surface $X$, which is the content of the quantitative main theorem, Theorem 1.5.
We proceed with the idea of the proof of Theorem 1.5. We set out to bound the following expression:
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}.$ (3.2)
We can rewrite the inner product as follows:
$|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})dt\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})h_{t}(\rho_{k})dt\right|}\,|\langle\psi_{j},a\psi_{k}\rangle|=\frac{\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}\,\psi_{k}\,dt\right\rangle\right|}{\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(\rho_{j})\,h_{t}(\rho_{k})\,dt\right|},$ (3.3)
where
@@ 3. Proof ideas
$T$ depends on $\delta$ via $T\delta=\tfrac{\pi}{2}$.
We refer to this as the geometric-spectral split, since the denominator contains the spectral data of the problem and the numerator contains the geometric data. Two different operators, $P_{t}$ and $P_{t,\tau}$, are used because we are dealing with off-diagonal averages where the separation of the eigenvalues is significant.
@@
The denominator of Expression (3.3) contains the spectral data. We acquire the following bounds for the denominator in Section 4:
The bound is obtained by carefully integrating and using the interrelations of $\delta,T,\rho_{j},$ and $\rho_{k}$. Since we are working in a compact energy window, $\rho_{j},\rho_{k}\leq\max(I)$. It is especially important that $\delta\in(0,\tfrac{2}{9}\min I)$. Being able to choose such a positive $\delta$ is crucial; this is why we have excluded the low-lying notes, i.e., those eigenfunctions of the Laplacian $\Delta_{X}$ with eigenvalues in the window $0,\tfrac{1}{4}]$.
Using the geometric-spectral split and the bounds for the spectral data gives an upper bound for Expression (3.2):
@@
$\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\psi_{k}\rangle\right|^{2}\leq 64\pi^{2}\max(I)^{4}\,\sum_{j:\rho_{j}\in I}\,\sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\left\langle\psi_{j},\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}dt\,\psi_{k}\right\rangle\right|^{2},$
where this has the following Hilbert-Schmidt norm upper bound by enlarging the double sum to include all basis elements:
$64\pi^{2}\max(I)^{4}\,\left\|\frac{1}{T}\int_{0}^{T}P_{t,\tau}^{*}aP_{t}dt\right\|_{HS}^{2}.$
The Hilbert-Schmidt norm contains the geometric data of the problem; it contains no eigenvalues or eigenfunctions and is now purely geometric.
In Section 5 we establish convenient forms for the integral kernels of $P_{t}$ and $P_{t,\tau}$ which are used to bound the Hilbert-Schmidt norm in Section 6. $P_{t}$ has integral kernel $K_{t}^{\Gamma}$:
$K_{t}^{\Gamma}(x,y)=\frac{1}{2\sqrt{2}\pi}\sum_{\gamma\in\Gamma}K_{t}(x,\gamma y),\qquad K_{t}(x,y)=\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}.$
$P_{t,\tau}$ has integral kernel $K_{t,\tau}^{\Gamma}=\cos(t\tau)K_{t}^{\Gamma}$. We notice that $P_{t}$ and $P_{t,\tau}$ are self-adjoint from these integral kernel formulas. Informally, the previous formulas for the integral kernels are obtained by using the inverse Selberg transform $\mathcal{S}^{-1}$ (e.g., see [LS17]):
$K_{t}^{\Gamma}(x,y)=\sum_{\gamma\in\Gamma}\mathcal{S}^{-1}(h_{t})(x,\gamma y).$
However, $h_{t}$ does not behave well enough for this to work immediately, and hence a more careful approach is needed; we approximate $P_{t}$ and $P_{t,\tau}$ to use analogous techniques.
By using the acquired formulas for the kernels, we bound the Hilbert-Schmidt norm in Section 6. Opening the Hilbert-Schmidt norm using the obtained kernel formulas and carefully rewriting it gives the following equivalent expression:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\cos(t\tau)\cos(t^{\prime}\tau)\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dzdt^{\prime}dt,$
where $D$ is a fundamental domain of $X$. The previous expression has the following upper bound obtained by taking absolute values, moving them inside, and bounding $|\cos(\cdot)|$ by $1$:
$\frac{1}{8\pi^{2}T^{2}}\int_{0}^{T}\int_{0}^{T}\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz\right|dt^{\prime}dt.$ (3.4)
Eliminating $\cos$ factors, which appear due to the simultaneous presence of $P_{t}$ and $P_{t,\tau}$, might lead to some loss of decay. A more careful approach could yield better estimates. $F_{t,t^{\prime},\rho}$ contains the core of the geometric data and is defined as follows:
$F_{t,t^{\prime},\rho}=\int_{B(z,t)\cap B(z^{\prime},t^{\prime})}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}.$
Here $z,z^{\prime}\in\mathbb{H}$ are any points separated by $\rho$. $F_{t,t^{\prime},\rho}$ is studied in Section 7. An especially important fact is that $F_{t,t^{\prime},\rho}=0$ when $\rho>t+t^{\prime}$.
In Expression (3.4), we split the integration over the fundamental domain $D$ into points with injectivity radius larger than $4T$, $D(4T)$, and into points with injectivity radius smaller than $4T$, $D\setminus D(4T)$. This approach leads to the following upper bound:
$\left|\int_{D}a(z)\int_{B(z,2T)}a(z^{\prime})^{*}F_{t,t^{\prime},d(z,z^{\prime})}^{2}dz^{\prime}dz\right|$ + $+\ 2\,\|a\|_{\infty}^{2}\int_{D\setminus D(4T)}\int_{B(z,2T)}F_{t,t^{\prime},d(z,z^{\prime})}\sum_{\gamma\in\Gamma}F_{t,t^{\prime},d(\gamma z,z^{\prime})}dz^{\prime}dz.$ (3.5)
For points with large injectivity radius, all group elements except $id\in\Gamma$ contribute zero, thus eliminating the sum. The latter term deals with points with small injectivity radius. The terms are treated separately.
The first term of Expression (3.5) can be written using polar coordinates around $z$ and has the following upper bound:
$\int_{0}^{2T}\sinh(\rho)F_{t,t^{\prime},\rho}^{2}\left|\langle a,a\circ\phi_{\rho}\rangle\right|d\rho\leq 30\|a\|_{2}^{2}\int_{0}^{2T}\sinh(\rho)(1+\rho)e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2},$
where the latter inequality is obtained using the exponential mixing theorem, Theorem 1.2. The exponential decay obtained via the exponential mixing theorem is perhaps the most crucial step in the entire proof. $\beta$ is a constant depending on the spectral gap.
The second term of Expression (3.5) is bounded more roughly using Hölder’s inequality. The interplay of $F_{t,t^{\prime},d(z,z^{\prime})}$ and $F_{t,t^{\prime},d(\gamma z,z^{\prime})}$ ensures that only finitely many group elements contribute. The upper bound can be written as:
@@
$8\pi^{2}\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{4T}}{\operatorname{InjRad}_{X}}\int_{0}^{2T}\sinh(\rho)(1+\rho)e^{-\beta\rho}F_{t,t^{\prime},\rho}^{2}.$
Using the above bounds, we obtain an upper bound for Expression (3.4):
$\frac{1}{T^{2}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|D\setminus D(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right)\int_{0}^{T}\int_{0}^{T}\int_{0}^{2T}\sinh(\rho)e^{-\beta\rho}(1+\rho)F_{t,t^{\prime},\rho}^{2}d\rho dt^{\prime}dt.$
This is straightforwardly bounded when we have the following bound for $F_{t,t^{\prime},\rho}$:
$F_{t,t^{\prime},\rho}\lesssim\sinh(\max(\rho,|t-t^{\prime}|))^{-\tfrac{1}{2}}.$ (3.6)
Obtaining this bound for $F_{t,t^{\prime},\rho}$ is the most technical part of the paper and is carried out in Section 7. The idea is to compute a weighted volume of the intersection of two hyperbolic balls using hyperbolic geometry and trigonometry. In practice, the integration is handled in multiple separate cases.
The final upper bound is a constant times
$\frac{1}{T\beta^{3}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ denotes the set of points with small injectivity radius. This shows that the double sum Expression (3.2) is bounded using the Hilbert-Schmidt norm bound:
$\max(I)^{4}\,\frac{1}{T\beta^{3}}\left(\|a\|_{2}^{2}+\|a\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right).$
A constant factor is omitted. This establishes the proof idea of Theorem 1.5.
We proceed to Theorems 1.3 and 1.4. We want to find an upper bound $\epsilon$ for a suitable $\delta$ for the left-hand side of Expression (3.1). Using the bounds from Theorem 1.5, we get the following upper bound (omitting a constant factor):
$\max(I)^{4}\frac{1}{T}\,\limsup_{n\to\infty}\frac{1}{N(X_{n},I)}\frac{1}{\beta(\lambda_{1}(X_{n}))^{3}}\left(\|a_{n}\|_{2}^{2}+\|a_{n}\|_{\infty}^{2}|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right).$ (3.7)
In the deterministic setting of Theorem 1.3, we use that $(a_{n})$ is uniformly bounded by $a_{max}$, $I$ is compact, (EXP) gives a universal lower bound for $(\beta(\lambda_{1}(X_{n})))$ denoted by $\beta_{min}$, and (UND) provides a uniform lower bound for $(\operatorname{InjRad}_{X_{n}})$ denoted by $\operatorname{InjRad}_{min}$. The upper bound is
$\frac{\max(I)^{4}a_{max}^{2}}{\beta_{min}^{3}}\frac{1}{T}\limsup_{n\to\infty}\left(\frac{|X_{n}|}{N(X_{n},I)}+\frac{|X_{n}\setminus X_{n}(4T)|\,e^{6T}}{N(X_{n},I)\,\operatorname{InjRad}_{min}}\right).$
The large-scale Weyl law gives $N(X_{n},I)\sim|X_{n}|$. Using Property (BSC), the volume ratio of points with small injectivity radius shrinks to zero. Hence, the upper bound becomes
$\frac{\max(I)^{4}a_{max}^{2}}{\beta_{min}^{3}}\frac{1}{T}.$
Since $T\delta=\tfrac{\pi}{2}$, we may choose $\delta$ sufficiently small so that the bound is less than $\epsilon$. This completes the proof idea for Theorem 1.3.
In the probabilistic setting of Theorem 1.4, a similar approach is taken; here $n=g$ is the genus of the surface. We do not have Properties (BSC), (EXP), (UND), or the large-scale Weyl law a priori. However, Theorem 8.2 provides analogous tools. For a fixed surface $X_{g}$ of genus $g$,
$\frac{1}{N(X_{g},I)\,\beta(\lambda_{1}^{(g)})^{3}}\left(|X_{g}|+|X_{g}\setminus X_{g}(4T)|\,\frac{e^{6T}}{\operatorname{InjRad}_{X_{g}}}\right)$
$\leq\ \frac{|X_{g}|}{|X_{g}|\,\mathcal{O}\left(\max(J)-\min(J)+\sqrt{\frac{\max(J)+1}{\log(g)}}\right)\,\left(1-\sqrt{1-4(\tfrac{1}{4}-\eta)}\right)^{3}}\,\left(1+\mathcal{O}(g^{-\frac{1}{4}})\,\frac{e^{6T}}{g^{-\frac{1}{24}}\,\log(g)^{\frac{9}{16}}}\right),$
with probability 1o(1)1-o(1), where $\eta\in(0,\tfrac{1}{4})$ and
$J=\left\{x^{2}+\tfrac{1}{4}\ :\ x\in I\right\}.$
Hence Expression (3.7) equals, with high probability in the high-genus limit (up to a constant factor),
$a_{max}^{2}\max(I)^{4}\frac{1}{T}.$
Since $T\delta=\tfrac{\pi}{2}$, we may choose $\delta$ sufficiently small so that the bound is less than $\epsilon$. This completes the proof idea for Theorem 1.4.